# Ammann Bars for Octagonal Tilings

### v5

@@ 1 Introduction
Shortly after the famous Penrose tilings were introduced by Roger Penrose in 1974 [23, 24, 25] and popularized by Martin Gardner in 1977 [15], amateur mathematician Robert Ammann [29] found particularly interesting decorations of the tiles (Figure 1): if one draws segments in the same way on all congruent tiles333taking into account the original markings on the sides that define the assembly rules. then on any valid tiling all those segments compose continuous straight lines, going in five different directions. Conversely if one follows the assembly rule consisting of prolonging every segment on the tiles into a straight line then the obtained tiling is indeed a Penrose tiling. Those lines are called Ammann bars and the corresponding matching rule is locally equivalent to the ones given by Penrose using arrows on the sides or alternative decorations.
@@
Penrose tilings have many interesting properties and can be generated in several ways. The cut and project method [^4] Terms in italics which are not defined in the introduction are defined formally in further sections. The introduction is meant to give a general idea of the article. follows their algebraic study by de Bruijn in 1981 [10]. Beenker soon proposed a whole family of tilings based on it [8], including the Ammann-Beenker tilings that Ammann found independently. A cut and project tiling can be seen as a discretization of a two-dimensional plane in an $n$-dimensional Euclidean space ($n>2$), and we will talk about $n\to 2$ tilings in that sense. When the slope of the plane does not contain any rational line, the tiling is non-periodic. This is the case for Penrose tilings for instance, so the set of tiles defining them is aperiodic: one can tile the plane with its tiles but only non-periodically. The first aperiodic tileset was found by Berger, thus proving the undecidability of the Domino Problem [9] and relating tilings to logic. Since then, relatively few others were exhibited: many non-periodic tilings exist (even infinitely many using the cut and project method), but we usually do not have a corresponding aperiodic tileset.
@@
Links were made between such tilings and quasicrystals [28, 32], that is crystals whose diffraction pattern is not periodic but still ordered, with rotational symmetries. The study of local rules, i.e. constraints on the way tiles can fit together in finite patterns, can help modeling the long range aperiodic order of quasicrystals. For instance, Penrose tilings are defined by their 1-atlas [12], which is a small number of small patterns: any and all tilings containing only those patterns (of the given size) are Penrose tilings. Alternately, they can also be defined by their Ammann local rules, as stated in the first paragraph. On the contrary, it was proven [11] that Ammann-Beenker tilings, also known as 8-fold tilings, do not have weak local rules (or “colored” local rules), i.e. no finite set of patterns (r-atlas) is enough to characterize them. Socolar found a variant of Ammann bars for them [30], but they extend outside the boundary of the tiles, thus do not fit the framework considered here.
@@
In addition to finding a suitable tileset, one of the difficulties lies in proving that this tileset, with only nearest-neighbor rules, can enforce projection tilings with arbitrary slopes. Goodman-Strauss [16] and then Fernique and Ollinger [13] designed constructions in the case of substitutive tilings: they showed that, under specific conditions, substitutive tilings have colored nearest-neighbor matching rules. In other words, substitutions can be forced by colored local rules. In the case of cut and project tilings, this is a corollary of a construction presented later by Fernique and Sablik [14] and based on works by Aubrun and Sablik [1], and Hochman [19]. This more general construction yields weak local rules for every computable slope, yet it is quite heavy, since it simulates Turing machines. The main result presented here is actually a corollary of Fernique and Sablik’s, but the construction is completely different and much lighter because it is specific to a narrower class of tilings, for (a priori) non-substitutive tilings, and additionally we exhibit more information on the structure of these tilings in the process. Specifically, we provide sufficient conditions for some cut and project tilings to have Ammann bars.
@@
Grünbaum and Shephard [17] detail the properties of Ammann bars in the case of Penrose tilings and their close relation to the Fibonacci word, which inspired Stehling to find an equivalent in 3-dimensional space [31]. They also present two tilesets by Ammann with Ammann bars (A2 and A3) but these are substitutive and not cut and project tilings. Generally speaking, many questions remain around Ammann bars and for now each family of aperiodic tilings has to be treated on a case-by-case basis, as Lück did in [22] with the examples of Ammann bars introduced by Socolar [30] and Ingalls [20]. Yet they can prove quite useful in the study of the structure of tilings, and were used by Porrier and Blondin Massé [26] to solve a combinatorial optimization problem on graphs defined by Penrose tilings.
@@
When it comes to $4\to 2$ tilings (discretizations [^5] A discretization of a 2-plane $P$ is, roughly speaking, a set of unit facets of $\mathbb{Z}^{4}$ which approximate the continuous plane $P$. The formal notion is planar, as defined in Section 2. of 2-dimensional planes in $\mathbb{R}^{4}$) and a few others like Penrose, which are $5\to 2$ tilings, the existence of weak local rules can be expressed in terms of subperiods, which are particular vectors of the slope [4, 6]. Careful examination of Penrose tilings from this angle shows that Ammann bars have the same directions as subperiods: there are two subperiods in each direction, one being $\varphi$ times longer than the other. Additionally, the lengths of the “integer versions” of subperiods are closely related to the distances between two consecutive Ammann bars in a given direction, as can be seen in Figure 1 (more details are given in Appendix A). Though interesting, this special case is too particular to hope for a generalization from it alone. Nonetheless, we think that Ammann bars are related to subperiods.
@@
Since subperiods are simpler for $4\to 2$ tilings, for which we also have a stronger result regarding weak local rules, we focus on those. Namely, Bédaride and Fernique [6] showed that a $4\to 2$ tiling has weak local rules if and only if its slope is characterized by its subperiods. As mentioned above, Ammann-Beenker tilings have no local rules and their slope cannot be characterized by its subperiods. The same holds for 12-fold tilings [3] for which Socolar also designed Ammann lines [30]. This means that these two types of tilings have decorated local rules, even though they have no weak local rules. In both cases, the “primary” Ammann bars found by Socolar have the same directions as projected subperiods but tiles decorated with those lines alone allow for periodic tilings of the plane, which is why Socolar introduced “secondary” Ammann bars. We conjecture that these lines, which are not related to subperiods, might force the correct frequencies of tiles, differentiating Ammann-Beenker tilings from other Beenker tilings, but we have not investigated further yet.
@@
We initially searched for examples of slopes based on two criteria: the slope has to be quadratic to meet the algebraic constraints (being characterized by subperiods), and we wanted “short” subperiods to enable human-eye observation of the corresponding drawings. The method involves a trial-and-error component: for each slope, we have to test whether it is possible to find a fine projection or not. Our first try was on Golden Octagonal tilings but their slope admits no fine projection. Hence we randomly generated potential candidates, found a few having a fine projection, and proved the above result on a particular example which we presented in the LATIN 2022 conference [27]. It consists of $4\to 2$ tilings defined by a given slope based on the irrationality of $\sqrt{3}$, that we called Cyrenaic tilings in reference to Theodorus of Cyrene who proved $\sqrt{3}$ to be irrational. In this case, the FP-method yields the set of decorated tiles depicted in Figure 2. Those tiles give Ammann bars to Cyrenaic tilings.
@@
The case of Penrose indicates that our construction could (and should) be adapted in order to work for $5\rightarrow 2$ tilings, or general cut and project ($n\to d$) tilings. In particular, for Penrose the lines are shifted and the number of lines is reduced compared with our method, so that only two decorated tiles are needed. Besides, in each direction the distance between two consecutive lines can take only two values, and the sequence of intervals is Sturmian and substitutive. In the case of Cyrenaic tilings, it is an open question whether the bi-infinite word defined by each sequence of intervals between Ammann bars is substitutive. If so, it may be that they could be composed according to the substitution. As for $n\to 3$ tilings, the case of icosahedral Ammann rhombohedral tilings could be studied as a first example of “Ammann planes” (instead of lines), since their slope is characterized by subperiods [5].
@@ 2.1 Canonical cut and project tilings
Let $e_{0},...,e_{n-1}$ be the canonical basis of $\mathbb{R}^{n}$. Following Levitov [21] and Bédaride and Fernique [4], a tiling by parallelograms can be lifted in $\mathbb{R}^{n}$, to correspond to a “stepped” surface of dimension 2 in $\mathbb{R}^{n}$, which is unique up to the choice of an initial vertex. An arbitrary vertex is first mapped onto the origin, then each tile of type $T_{ij}$ is mapped onto the 2-dimensional face of a unit hypercube of $\mathbb{Z}^{n}$ generated by $e_{i}$ and $e_{j}$, such that two tiles adjacent along an edge $v_{i}$ are mapped onto two faces adjacent along an edge $e_{i}$. This is particularly intuitive for $3\to 2$ tilings which are naturally seen in $3$ dimensions (Fig. 3, left). The principle is the same for larger $n$, though difficult to visualize.
@@
Strongly planar tilings by parallelograms can also be obtained by the (canonical) cut and project method. For this, consider a $d$-dimensional affine plane $E\subset\mathbb{R}^{n}$ such that $E\cap\mathbb{Z}^{n}=\emptyset$, select (“cut”) all the $d$-dimensional facets of $\mathbb{Z}^{n}$ which lie within the tube $E+[0,1]^{n}$, then “project” them onto $\mathbb{R}^{d}$. If this projection $\pi$ yields a tiling of $\mathbb{R}^{d}$ it is called valid (see Figure 4), and the tiling is a strongly planar tiling by parallelograms with slope $E$. Such tilings are called canonical cut and project tilings or simply n $\to$ d tilings. Not every projection is suitable, but the orthogonal projection onto $E$ seen as $\mathbb{R}^{d}$ is known to be valid [18]. Here we only consider the case of a $2$-dimensional slope $E$ which is totally irrational, that is, which does not contain any rational line. This yields aperiodic tilings of the plane.
@@
where $\alpha\approx 1.89$ is the only real root of $x^{3}-x^{2}-x-1$. Ammann-Beenker tilings, composed of tiles of the set A5 in the terminology of Grünbaum and Shephard [17], are the $4\rightarrow 2$ tilings with slope $E$ generated by
@@
where $\varphi=(1+\sqrt{5})/2$ is the golden ratio. The “strict” Penrose tilings as defined by Roger Penrose in [24] (set P3 in the terminology of [17]) correspond to the case when $E$ contains a point whose coordinates sum to an integer.
@@ 2.2 Local rules
Firstly, weak local rules for a tiling $T$ can be defined as in [4]. A pattern is a connected finite subset of tiles of $T$. Following [21], an r-map of $T$ is a pattern formed by the tiles of $T$ which intersect a closed disk of radius $r\geq 0$. The r-atlas of $T$, denoted by $T(r)$, is then the set of all $r$-maps of $T$ (up to translation). In the case of a canonical cut and project tiling, it is a finite set. A canonical cut and project tiling $\mathcal{P}$ of slope $E$ is said to admit weak local rules if there exist $r\geq 0$ and $t\geq 1$, respectively called radius and thickness, such that any $n\to d$ tiling $T$ whose $r$-atlas is contained in $\mathcal{P}(r)$ is planar with slope $E$ and thickness at most $t$. By extension, the slope $E$ is then said to admit local rules. In that case, we say that the slope of $\mathcal{P}$ is characterized by its patterns of a given size. Local rules are strong if $t=1$. Penrose tilings have strong local rules and the slope is characterized by patterns of the 1-atlas if the sides of the tiles have length 1 (see [28], Theorem 6.1, p.177).
@@ 2.3 Subperiods
Adapted from Bédaride and Fernique [2], the $i_{1},...,i_{n-3}$-shadow of an $n\to 2$ tiling $T$ is the orthogonal projection $\omega_{i_{1},...,i_{n-3}}$ of its lift on the space generated by $\{e_{j}\mid 0\leq j\leq n-1,j\neq i_{1},...,i_{n-3}\}$. This corresponds to reducing to zero the lengths of $\pi(e_{i_{1}}),...,\pi(e_{i_{n-3}})$ in the tiling, so that the tiles defined by these vectors disappear. This is illustrated in Figure 5. An $n\to 2$ tiling thus has ($n\choose 3$ shadows.
@@
This notion was first introduced by Levitov [21] as the second intersection condition and then developed by Bédaride and Fernique, who showed in [4] and [6] that in the case of $4\rightarrow 2$ tilings, a plane admits weak local rules if and only if it is determined by its subperiods. It was shown in [2] that this is not the case for Ammann-Beenker tilings: indeed, their subperiods are also subperiods of all Beenker tilings (introduced in [8]), that are the planar tilings with a slope generated, for any $s\in(0,\infty)$, by
@@
The Ammann-Beenker tilings correspond to the case $s=\sqrt{2}$ and do not admit local rules. On the other hand, generalized Penrose tilings have a slope characterized by its subperiods [4] and do admit local rules.

### v6

@@ Abstract
keywords:
aperiodicity, tilings, symbolic dynamics, local rules
@@ 1 Introduction
Shortly after the famous Penrose tilings were introduced by Roger Penrose (1974, 1978, 1979) and popularized by Martin Gardner (1977), amateur mathematician Robert Ammann (Senechal, 2004) found particularly interesting decorations of the tiles (Figure 1): if one draws segments in the same way on all congruent tiles111taking into account the original markings on the sides that define the assembly rules. then on any valid tiling all those segments compose continuous straight lines, going in five different directions. Conversely if one follows the assembly rule consisting of prolonging every segment on the tiles into a straight line then the obtained tiling is indeed a Penrose tiling. Those lines are called Ammann bars and the corresponding matching rule is locally equivalent to the ones given by Penrose using arrows on the sides or alternative decorations.
@@
Penrose tilings have many interesting properties and can be generated in several ways. The cut and project method [^2] Terms in italics which are not defined in the introduction are defined formally in further sections. The introduction is meant to give a general idea of the article. follows their algebraic study by de Bruijn (1981). Beenker (1982) soon proposed a whole family of tilings based on it , including the Ammann-Beenker tilings that Ammann found independently. A cut and project tiling can be seen as a discretization of a two-dimensional plane in an $n$-dimensional Euclidean space ($n>2$), and we will talk about $n\to 2$ tilings in that sense. When the slope of the plane does not contain any rational line, the tiling is non-periodic. This is the case for Penrose tilings for instance, so the set of tiles defining them is aperiodic: one can tile the plane with its tiles but only non-periodically. The first aperiodic tileset was found by Berger, thus proving the undecidability of the Domino Problem (Berger, 1966) and relating tilings to logic. Since then, relatively few others were exhibited: many non-periodic tilings exist (even infinitely many using the cut and project method), but we usually do not have a corresponding aperiodic tileset.
@@
Links were made between such tilings and quasicrystals (Senechal, 1995; Steinhardt and Ostlund, 1987), that is crystals whose diffraction pattern is not periodic but still ordered, with rotational symmetries. The study of local rules, i.e. constraints on the way tiles can fit together in finite patterns, can help modeling the long range aperiodic order of quasicrystals. For instance, Penrose tilings are defined by their 1-atlas (Fernique and Lutfalla, 2023), which is a small number of small patterns: any and all tilings containing only those patterns (of the given size) are Penrose tilings. Alternately, they can also be defined by their Ammann local rules, as stated in the first paragraph. On the contrary, it was proven (Bédaride and Fernique, 2015b) that Ammann-Beenker tilings, also known as 8-fold tilings, do not have weak local rules (or “colored” local rules), i.e. no finite set of patterns (r-atlas) is enough to characterize them. Socolar (1989) found a variant of Ammann bars for them, but they extend outside the boundary of the tiles, thus do not fit the framework considered here.
@@
In addition to finding a suitable tileset, one of the difficulties lies in proving that this tileset, with only nearest-neighbor rules, can enforce projection tilings with arbitrary slopes. Goodman-Strauss (1998) and then Fernique and Ollinger (2010) designed constructions in the case of substitutive tilings: they showed that, under specific conditions, substitutive tilings have colored nearest-neighbor matching rules. In other words, substitutions can be forced by colored local rules. In the case of cut and project tilings, this is a corollary of a construction presented later by Fernique and Sablik (2012) and based on works by Aubrun and Sablik (2013), and Hochman (2009). This more general construction yields weak local rules for every computable slope, yet it is quite heavy, since it simulates Turing machines. The main result presented here is actually a corollary of Fernique and Sablik’s, but the construction is completely different and much lighter because it is specific to a narrower class of tilings, for (a priori) non-substitutive tilings, and additionally we exhibit more information on the structure of these tilings in the process. Specifically, we provide sufficient conditions for some cut and project tilings to have Ammann bars.
@@
Grünbaum and Shephard (1987) detail the properties of Ammann bars in the case of Penrose tilings and their close relation to the Fibonacci word, which inspired Stehling to find an equivalent in 3-dimensional space (Stehling, 1992). They also present two tilesets by Ammann with Ammann bars (A2 and A3) but these are substitutive and not cut and project tilings. Generally speaking, many questions remain around Ammann bars and for now each family of aperiodic tilings has to be treated on a case-by-case basis, as Lück (1993) did with the examples of Ammann bars introduced by Socolar (1989) and Ingalls (1993). Yet they can prove quite useful in the study of the structure of tilings, and were used by Porrier and Blondin Massé (2020) to solve a combinatorial optimization problem on graphs defined by Penrose tilings.
@@
When it comes to $4\to 2$ tilings (discretizations [^3] A discretization of a 2-plane $P$ is, roughly speaking, a set of unit facets of $\mathbb{Z}^{4}$ which approximate the continuous plane $P$. The formal notion is planar, as defined in Section 2. of 2-dimensional planes in $\mathbb{R}^{4}$) and a few others like Penrose, which are $5\to 2$ tilings, the existence of weak local rules can be expressed in terms of subperiods, which are particular vectors of the slope (Bédaride and Fernique, 2015a, 2017). Careful examination of Penrose tilings from this angle shows that Ammann bars have the same directions as subperiods: there are two subperiods in each direction, one being $\varphi$ times longer than the other. Additionally, the lengths of the “integer versions” of subperiods are closely related to the distances between two consecutive Ammann bars in a given direction, as can be seen in Figure 1 (more details are given in Appendix A). Though interesting, this special case is too particular to hope for a generalization from it alone. Nonetheless, we think that Ammann bars are related to subperiods.
@@
Since subperiods are simpler for $4\to 2$ tilings, for which we also have a stronger result regarding weak local rules, we focus on those. Namely, Bédaride and Fernique (2017) showed that a $4\to 2$ tiling has weak local rules if and only if its slope is characterized by its subperiods. As mentioned above, Ammann-Beenker tilings have no local rules and their slope cannot be characterized by its subperiods. The same holds for 12-fold tilings (Bédaride and Fernique, 2015b) for which Socolar also designed Ammann lines (Socolar, 1989). This means that these two types of tilings have decorated local rules, even though they have no weak local rules. In both cases, the “primary” Ammann bars found by Socolar have the same directions as projected subperiods but tiles decorated with those lines alone allow for periodic tilings of the plane, which is why Socolar introduced “secondary” Ammann bars. We conjecture that these lines, which are not related to subperiods, might force the correct frequencies of tiles, differentiating Ammann-Beenker tilings from other Beenker tilings, but we have not investigated further yet.
@@
We initially searched for examples of slopes based on two criteria: the slope has to be quadratic to meet the algebraic constraints (being characterized by subperiods), and we wanted “short” subperiods to enable human-eye observation of the corresponding drawings. The method involves a trial-and-error component: for each slope, we have to test whether it is possible to find a fine projection or not. Our first try was on Golden Octagonal tilings but their slope admits no fine projection. Hence we randomly generated potential candidates, found a few having a fine projection, and proved the above result on a particular example which we presented in the LATIN 2022 conference (Porrier and Fernique, 2022). It consists of $4\to 2$ tilings defined by a given slope based on the irrationality of $\sqrt{3}$, that we called Cyrenaic tilings in reference to Theodorus of Cyrene who proved $\sqrt{3}$ to be irrational. In this case, the FP-method yields the set of decorated tiles depicted in Figure 2. Those tiles give Ammann bars to Cyrenaic tilings.
@@
The case of Penrose indicates that our construction could (and should) be adapted in order to work for $5\rightarrow 2$ tilings, or general cut and project ($n\to d$) tilings. In particular, for Penrose the lines are shifted and the number of lines is reduced compared with our method, so that only two decorated tiles are needed. Besides, in each direction the distance between two consecutive lines can take only two values, and the sequence of intervals is Sturmian and substitutive. In the case of Cyrenaic tilings, it is an open question whether the bi-infinite word defined by each sequence of intervals between Ammann bars is substitutive. If so, it may be that they could be composed according to the substitution. As for $n\to 3$ tilings, the case of icosahedral Ammann rhombohedral tilings could be studied as a first example of “Ammann planes” (instead of lines), since their slope is characterized by subperiods (Bédaride and Fernique, 2016).
@@ 2.1 Canonical cut and project tilings
Let $e_{0},...,e_{n-1}$ be the canonical basis of $\mathbb{R}^{n}$. Following Levitov (1988) and Bédaride and Fernique (2015a), a tiling by parallelograms can be lifted in $\mathbb{R}^{n}$, to correspond to a “stepped” surface of dimension 2 in $\mathbb{R}^{n}$, which is unique up to the choice of an initial vertex. An arbitrary vertex is first mapped onto the origin, then each tile of type $T_{ij}$ is mapped onto the 2-dimensional face of a unit hypercube of $\mathbb{Z}^{n}$ generated by $e_{i}$ and $e_{j}$, such that two tiles adjacent along an edge $v_{i}$ are mapped onto two faces adjacent along an edge $e_{i}$. This is particularly intuitive for $3\to 2$ tilings which are naturally seen in $3$ dimensions (Fig. 3, left). The principle is the same for larger $n$, though difficult to visualize.
@@
Strongly planar tilings by parallelograms can also be obtained by the (canonical) cut and project method. For this, consider a $d$-dimensional affine plane $E\subset\mathbb{R}^{n}$ such that $E\cap\mathbb{Z}^{n}=\emptyset$, select (“cut”) all the $d$-dimensional facets of $\mathbb{Z}^{n}$ which lie within the tube $E+[0,1]^{n}$, then “project” them onto $\mathbb{R}^{d}$. If this projection $\pi$ yields a tiling of $\mathbb{R}^{d}$ it is called valid (see Figure 4), and the tiling is a strongly planar tiling by parallelograms with slope $E$. Such tilings are called canonical cut and project tilings or simply n $\to$ d tilings. Not every projection is suitable, but the orthogonal projection onto $E$ seen as $\mathbb{R}^{d}$ is known to be valid (Harriss, 2004). Here we only consider the case of a $2$-dimensional slope $E$ which is totally irrational, that is, which does not contain any rational line. This yields aperiodic tilings of the plane.
@@
where $\alpha\approx 1.89$ is the only real root of $x^{3}-x^{2}-x-1$. Ammann-Beenker tilings, composed of tiles of the set A5 in the terminology of Grünbaum and Shephard (1987), are the $4\rightarrow 2$ tilings with slope $E$ generated by
@@
where $\varphi=(1+\sqrt{5})/2$ is the golden ratio. The “strict” Penrose tilings as defined by Roger Penrose (1978) (set P3 in the terminology of Grünbaum and Shephard (1987)) correspond to the case when $E$ contains a point whose coordinates sum to an integer.
@@ 2.2 Local rules
Firstly, weak local rules for a tiling $T$ can be defined as in Bédaride and Fernique (2015a). A pattern is a connected finite subset of tiles of $T$. Following Levitov (1988), an r-map of $T$ is a pattern formed by the tiles of $T$ which intersect a closed disk of radius $r\geq 0$. The r-atlas of $T$, denoted by $T(r)$, is then the set of all $r$-maps of $T$ (up to translation). In the case of a canonical cut and project tiling, it is a finite set. A canonical cut and project tiling $\mathcal{P}$ of slope $E$ is said to admit weak local rules if there exist $r\geq 0$ and $t\geq 1$, respectively called radius and thickness, such that any $n\to d$ tiling $T$ whose $r$-atlas is contained in $\mathcal{P}(r)$ is planar with slope $E$ and thickness at most $t$. By extension, the slope $E$ is then said to admit local rules. In that case, we say that the slope of $\mathcal{P}$ is characterized by its patterns of a given size. Local rules are strong if $t=1$. Penrose tilings have strong local rules and the slope is characterized by patterns of the 1-atlas if the sides of the tiles have length 1 (see Senechal (1995), Theorem 6.1, p.177).
@@ 2.3 Subperiods
Adapted from Bédaride and Fernique (2013), the $i_{1},...,i_{n-3}$-shadow of an $n\to 2$ tiling $T$ is the orthogonal projection $\omega_{i_{1},...,i_{n-3}}$ of its lift on the space generated by $\{e_{j}\mid 0\leq j\leq n-1,j\neq i_{1},...,i_{n-3}\}$. This corresponds to reducing to zero the lengths of $\pi(e_{i_{1}}),...,\pi(e_{i_{n-3}})$ in the tiling, so that the tiles defined by these vectors disappear. This is illustrated in Figure 5. An $n\to 2$ tiling thus has ($n\choose 3$ shadows.
@@
This notion was first introduced by Levitov (1988) as the second intersection condition and then developed by Bédaride and Fernique (2015a, 2017), who showed that in the case of $4\rightarrow 2$ tilings, a plane admits weak local rules if and only if it is determined by its subperiods. It was shown by Bédaride and Fernique (2013) that this is not the case for Ammann-Beenker tilings: indeed, their subperiods are also subperiods of all Beenker tilings (introduced by Beenker (1982)), that are the planar tilings with a slope generated, for any $s\in(0,\infty)$, by
@@
The Ammann-Beenker tilings correspond to the case $s=\sqrt{2}$ and do not admit local rules. On the other hand, generalized Penrose tilings have a slope characterized by its subperiods (Bédaride and Fernique, 2015a) and do admit local rules.